OpenStudy (anonymous):

A line has a vertical change per unit of horizontal change of -5/4 and passes through point ( 2, 2) . Which is another point on the line? 1. (-3, 6) 2. ( -10, 8) 3. ( 8, -10) 4. (6, -3)

OpenStudy (anonymous):

u have dy/dx just use y=dy/dx +c and get c by satisfying the given point

OpenStudy (anonymous):

find the equation using point slope form \[y-y _{1}=m(x-x _{1})\] where m is the slope

OpenStudy (ash2326):

@Olympian2016 Do you know what is meant by "vertical change over horizontal change " ?

OpenStudy (anonymous):

it is dy/dx or just slope ash2326

OpenStudy (anonymous):

can someone give me a step by step process to get the answer

OpenStudy (anonymous):

y=-5/4x+c now satisfy point (2,2) and get c then satisfy ur options and the one which agrees with ur eqation is the anwer

OpenStudy (anonymous):

so the answer is -3,6